Atsie (atsie):

I'm confused on part b) and c) If its a word problem, then is it supposed to be written that way? Or is it a story problem where Mary had this many apples or Jack had this many apples? If part b) is right, then how on earth do I answer part c) ? (pic below)

Atsie (atsie):



OpenStudy (anonymous):

Where

OpenStudy (anonymous):

oh

Atsie (atsie):

@skullpatrol @Agl202

OpenStudy (texaschic101):

cant see the image very good...too blurry

Atsie (atsie):

Here, view the attachment. :)

Atsie (atsie):

@texaschic101 :)

OpenStudy (texaschic101):

1/2 (x - 4) = 1/3x + 2....multiply both sides by 6 3(x - 4) = 2x + 12...distribute thru parenthesis 3x - 12 = 2x + 12...subtract 2x...add 12 3x - 2x = 12 + 12 x = 24

OpenStudy (texaschic101):

part b...i think they want a story problem

Atsie (atsie):

Okay, one minute

OpenStudy (texaschic101):

a little off...if jane gave her apples away, she would have none left

OpenStudy (texaschic101):

jane and mary started with the same # of apples. jane had 4 rotten appples..so she threw them away. she then gave 1/2 of her apples away. mary gave 1/3 of her apples, along with the 2 she found, away. how many apples did they start with ? kinda tupid...but i cant think of a good story right now

Atsie (atsie):

Okay, so using your story problem, how do we answer part C) ?

OpenStudy (texaschic101):

already did that. worked it out and x = 24...so if u go with the story, jane and mary both started with 24 apples

OpenStudy (texaschic101):

i worked it out ^^ up there a few posts

Atsie (atsie):

Ooh, I half missed that part. Alright, I understand what I'm doing now. Half distracted with English. Me and two subjects don't get along. Thank you so much for your help. I truly appreciate it!! ^.^

OpenStudy (texaschic101):

lol...no prob